OpenStudy (anonymous):

I have been trying to figure this out for awhile with no luck. can anyone help? suppose that a particles position is described by r(t)=(t^2+1)i+(t-5)j Give an equation(in the form of a formula involving x and y set equal to 0)whose solutions consist of the path of the particle. The answer looks like ____________=0

OpenStudy (anonymous):

\[ y=t-5\\ t=y+5\\ x=t^2+1=(y+5)^2 +1\\ x-(y+5)^2 -1=0 \]

OpenStudy (anonymous):

You have to eliminate t between x and y

OpenStudy (anonymous):

ohhh ok so you solved for t on one of them and plugged it into the other. thats it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright cool tyvm

OpenStudy (anonymous):

yw